# DynST: Dynamic Sparse Training for Resource-Constrained Spatio-Temporal Forecasting

### v1

@@ 6.4. Structural & Ablation study ($\mathcal{RQ}$4)
Table 3. SSIM and PSNR results on three research domain. The underline symbol represents the best performance. Ori denotes the original results, +Dyn denotes add DynST at sparsity 30%.
@@
Earthfarseer (TaxiBJ+) $\underline{0.98}$${}_{\pm 0.01}$/ $\underline{0.96}$${}_{\pm 0.01}$ $\underline{37.84}$${}_{\pm 0.01}$/ $\underline{36.44}$${}_{\pm 0.01}$
@@
Earthfarseer (WeatherBench) $\underline{0.98}$±${}_{\pm 0.01}$/ $\underline{0.97}$${}_{\pm 0.01}$ $\underline{39.27}$${}_{\pm 0.11}$/ $\underline{38.12}$${}_{\pm 0.03}$
@@
Earthfarseer (FIT) $\underline{0.95}$${}_{\pm 0.01}$ / $\underline{0.93}$${}_{\pm 0.01}$ $\underline{37.23}$${}_{\pm 0.01}$/ $\underline{36.31}$${}_{\pm 0.01}$
@@
In the last, we select three training schemes (Earthfarseer as the base model) to explore the performance of our algorithm and the benefits of combining our algorithm with mainstream training approaches: (1) One-shot pruning (OP): We thoroughly train our model and subsequently conduct training of the mask for a one-time pruning process. (2) Iterative pruning (IP): As our work can be regarded as a pruning method, we have opted for the widely recognized iterative pruning (IP) strategy (Frankle and Carbin, 2018) in the main manuscript part, we prune data for 10 times and every time for pruning 4% sub-counterpart. (3) Dynamic sparse training (DST): We select a target sparsity level and then maintain the data training consistently at this fixed sparsity. Dynamically, we remove and restore the smallest and largest magnitudes in the mask (Evci et al., 2020).We set a 40% sparsity for dynamic training. Table 4 shows STGCN, SimVP, and Earthfarseer’s performance in IP, OS, and DST training methods. Their RMSEs are 0.5698, 0.5108, 0.3507 (IP), 0.6197, 0.5650, 0.4121 (OS), and 0.5792, 0.5261, 0.3495 (DST).
@@ 7. Conclusion
In this paper, we introduce the concept of dynamic sparse training in the context of sensor deployment, termed DynST, which adjusts sensor deployment dynamically through training without compromising the model’s predictive capabilities. DynST ingeniously circumvents the complexity issues posed by the temporal dimension through clever dimension mapping. Following this, through dynamic training and mask operations, we can precisely identify the less significant parts of the output data, which correspond to the areas detected by the sensors. DynST is both general and succinct, compatible with many mainstream training schemes, such as one-shot, iterative pruning and dynamical sparse training, to boost inference speed without significant performance degradation.
@@ Acknowledgements.
This work is also supported by Guangzhou-HKUST (GZ) Joint Funding Program (No. 2024A03J0620)
@@
Appendix C Algorithm
Algorithm 1 Dynamic Sparse Training (DynST) Framework
0:  Input graph $\mathcal{G}_{in}$, Network $f$, Target Sparsity $S_{g}\%$
0:  Sparse mask $M_{g}$
1:  Initialize graph mask $M_{g}$
3:  while $1-\frac{\|M_{g}\|_{0}}{\|\mathcal{G}_{in}[;*]\|_{0}}<S_{g}$ do
4:     Training network for $R$ iterations
5:     Training $M_{g}$ for $M$ iterations
6:     Dynamical sparse training using Eq.5 and Eq.6
7:     Adjust $M_{g}$ using Eq.7
8:  end while
9:  $\hat{\mathcal{G}}_{in}\leftarrow\hat{\mathcal{G}}_{in}\odot M_{g}$

### v2

@@ 6.4. Structural & Ablation study ($\mathcal{RQ}$4)
Table 3. SSIM and PSNR results on three research domains. The underline symbol represents the best performance. Ori denotes the original results, +Dyn denotes add DynST at sparsity 30%.
@@
Earthfarseer (TaxiBJ+) $\underline{0.98}$0.01/ $\underline{0.96}$0.01 $\underline{37.84}$0.01/ $\underline{36.44}$0.01
@@
Earthfarseer (WeatherBench) $\underline{0.98}$±0.01/ $\underline{0.97}$0.01 $\underline{39.27}$0.11/ $\underline{38.12}$0.03
@@
Earthfarseer (FIT) $\underline{0.95}$0.01 / $\underline{0.93}$0.01 $\underline{37.23}$0.01/ $\underline{36.31}$0.01
@@
In the last, we select three training schemes (Earthfarseer as the base model) to explore the performance of our algorithm and the benefits of combining our algorithm with mainstream training approaches: (1) One-shot pruning (OP): We thoroughly train our model and subsequently conduct training of the mask for a one-time pruning process. (2) Iterative pruning (IP): As our work can be regarded as a pruning method, we have opted for the widely recognized iterative pruning (IP) strategy (Frankle and Carbin, 2018) in the main manuscript part, we prune data for 10 times and every time for pruning 4% sub-counterpart. (3) Dynamic sparse training (DST): We select a target sparsity level and then maintain the data training consistently at this fixed sparsity. Dynamically, we remove and restore the smallest and largest magnitudes in the mask (Evci et al., 2020). We set a 40% sparsity for dynamic training. Table 4 shows STGCN, SimVP, and Earthfarseer’s performance in IP, OS, and DST training methods. Their RMSEs are 0.5698, 0.5108, 0.3507 (IP), 0.6197, 0.5650, 0.4121 (OS), and 0.5792, 0.5261, 0.3495 (DST). In summary, our DST performs the best.
@@ 7. Conclusion
In this paper, we introduce the concept of dynamic sparse training in the context of sensor deployment, termed DynST, which adjusts sensor deployment dynamically through training without compromising the model’s predictive capabilities. DynST ingeniously circumvents the complexity issues posed by the temporal dimension through clever dimension mapping. Following this, through dynamic training and mask operations, we can precisely identify the less significant parts of the output data, which correspond to the areas detected by the sensors.
@@ Acknowledgements.
This work is mainly supported by the National Natural Science Foundation of China (No. 62402414). This work is also supported by Tencent (CCF-Tencent Open Fund, Tencent Rhino-Bird Focused Research Program), Didi (CCF-DiDi GAIA Collaborative Research Funds), Guangzhou Municipal Science and Technology Project (No. 2023A03J0011) and Guangzhou-HKUST(GZ) Joint Funding Program (No. 2024A03J0620).